A truck is towing a 1200 kg car. The rope being used can only with stand a tension of 800 N a. How much work does the truck do in towing the car 20m? b. What is the maximum acceleration of the truck before the rope breaks?
Answer #1

a)

W = Fd = 800 * 20 = 16000 J

=====

b)

a = F/m = 800 / 1200 = 0.667 m/s^2

=====

Comment in case any doubt, will reply for sure.. Goodluck